On January 1, 2018, Majestic Mantles leased a lathe from Equipment Leasing under a finance lease. Lease payments are made annually. Title does not transfer to the lessee and there is no purchase option or guarantee of a residual value by Majestic. Portions of the Equipment Leasing’s lease amortization schedule appear below: Jan. 1 Payments Effective Interest Decrease in Balance Outstanding Balance 2018 280,948 2018 30,000 30,000 250,948 2019 30,000 25,095 4,905 246,043 2020 30,000 24,604 5,396 240,647 2021 30,000 24,065 5,935 234,712 2022 30,000 23,471 6,529 228,183 2023 30,000 22,818 7,182 221,001 2024 30,000 22,100 7,900 213,101 — — — — — — — — — — — — — — — 2035 30,000 7,461 22,539 52,068 2036 30,000 5,207 24,793 27,275 2037 30,000 2,727 27,275 0 Required: 1. What is Majestic’s lease liability at the beginning of the lease (after the first payment)? 2. What amount would Majestic record as a right-of-use asset? 3. What is the lease term in years? 4. What is the effective annual interest rate? (Round your percentage answers to 1 decimal place.) 5. What is the total amount of lease payments? 6. What is the total effective interest expense recorded over the term of the lease?
Answer #1

Part 1

Majestic’s lease liability at the beginning of the lease (after the first payment) = 280948-30000= $250948

Part 2

Right-of-use asset = $280948

Part 3

Lease term = 20 years (Jan 1 2018 - Jan 1 2037)

Part 4

Effective interest rate = 25095/250948 = 10.0%

Part 5

Total amount of lease payments = annual payment*lease years = 30000*20 = $600000

Part 6

Total effective interest expense = total amount of lease payments-beginning lease liability= 600000-280948 = $319052